# Limits of Probability Measures with General Coefficients

### v2

@@ 6. Polynomial Coefficients
When $[1]\mathcal{D}^{\theta}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))$ is expressed as a polynomial in the $c_{F}^{\nu}$ for $\nu\in P^{+}$, we expect the coefficients to be polynomials in $\theta$ and $N$. In this section, we show that this is the case and relate the leading terms of the polynomial coefficients with the main results. In fact, we use this characterization of the coefficients to prove Theorem 1.9, which generalizes Theorem 1.7.
6.1. Proof of Theorem 1.9
First, we state the following result, which is a straightforward implication of Lemma 4.7.
@@
    If $\sigma(\mathfrak{r}_{1})=\sigma(\mathfrak{r}_{2})$, then $[1]\mathcal{D}_{\mathfrak{r}_{2}}^{\theta}(F(x_{1},\ldots,x_{N}))=[1]\mathcal{D}_{\mathfrak{r}_{1}}^{\theta}(F(x_{1},\ldots,x_{N}))$ for $N\geq\max(\mathfrak{r}_{1},\mathfrak{r}_{2})$.
@@
Suppose $\mathfrak{r}_{1}=\{i_{j}^{1}\}_{1\leq j\leq k}$ and $\mathfrak{r}_{2}=\{i_{j}^{2}\}_{1\leq j\leq k}$, where $k\geq 1$. Clearly, there is nothing to prove if $\mathfrak{r}_{1}$ and $\mathfrak{r}_{2}$ have different lengths. Statement (a) follows from Lemma 4.7, because
@@
there exists a polynomial $f(x,y)$ with rational coefficients such that for all complex numbers $\theta$ and $N\geq\max(\mathfrak{r})$, the coefficient of $p$ in $[1]\mathcal{D}^{\theta}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))$ is $f(\theta,N)$.
@@
Let $T_{N}$ be the set of $s\in U_{N}$ such that $P(x_{1},\ldots,x_{N})$ is nonzero and has degree 00 in the $x_{i}$ and $\prod_{j=1}^{r}c_{F}^{\nu_{j}^{\prime}}=p$. If $s\in T_{N}$ and $r(s(\theta))_{k}=d\theta^{l(s)}p$ for $\0d\in\mathbb{R}\backslash\{0\}$, let $C^{\prime}(s)=d$; note that $C^{\prime}(s)$ does not depend on $\theta$. The coefficient of $p$ in $[1]\mathcal{D}^{\theta}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))$ is $\sum_{s\in T_{N}}C^{\prime}(s)\theta^{l(s)}$ for all $\theta$.
If the sum of the $|\nu_{i}|$ is not $k$, from Corollary 3.6 the coefficient of $p$ is 00. Suppose that the sum of the $|\nu_{i}|$ is $k$. Let $S$ be the set of elements in $\mathfrak{r}$. Also, for $s\in T_{N}$, let $X(s)$ be the set $x_{i}$ such that $i\notin\mathfrak{r}$ and $x_{i}$ appears in switches or term multiplications of $s$. Since each $x_{i}\in X(s)$ from term multiplications must be eliminated by at least one switch by 3.10.1, and each other $x_{i}\in X(s)$ appears in at least one switch, $|X(s)|\leq k$ because there are at most $k$ switches.
For each $A\subset\{x_{1},\ldots,x_{N}\}\backslash S$, where the number of $s\in T_{N}$ such that $X(s)=A$ is finite, let
@@
For $A$ with $|A|=Q$ to exist, we must have $N\geq|S|+Q$. Suppose $Q$ is an integer, $0\leq Q\leq k$. By symmetry, $g(A)$ is the same for all $A\subset\{x_{1},\ldots,x_{N}\}\backslash S$, such that $|A|=Q$. With this, let
@@
Using Lemma 6.2, for indices $\mathfrak{r}$ of length $k$ and $N\geq\max(\mathfrak{r})$, let the coefficient of $p$ in $[1]\mathcal{D}^{\theta}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))$ be $f_{p,\mathfrak{r}}(\theta,N)$ for $p$ in $\mathfrak{P}_{k}$. Then, for $N\geq\max(\mathfrak{r})$,
(23) $[1]\mathcal{D}^{\theta}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))=\sum_{p\in\mathfrak{P}_{k}}f_{p,\mathfrak{r}}(\theta,N)p.$
Also, from Proposition 3.10 with $c=0$ and $\theta_{N}=\theta$ for $N\geq 1$ for any $\theta\in\mathbb{C}$, $f_{p,\mathfrak{r}}(x,y)$ has degree at most $k-\ell(p)$ in $y$. Then, let the coefficient of $y^{k-\ell(p)}$ in $f_{p,\mathfrak{r}}(x,y)$ be $s_{p,\mathfrak{r}}(x)$, with $s_{p,\mathfrak{r}}(x)=0$ being possible.
@@
Proof of Lemma 6.3.
Consider the indices $\mathfrak{r}^{\prime}$ of length $k$, where the first $\lambda_{1}$ indices are $1$, the next $\lambda_{2}$ are $2$, and so forth, until the last $\lambda_{m}$ indices are $m$. From Corollary 6.1, $[1]\mathcal{D}^{\theta}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))=[1]\mathcal{D}^{\theta}_{\mathfrak{r}^{\prime}}(F(x_{1},\ldots,x_{N}))$ for all $N\geq\max(\mathfrak{r})$. Therefore, $f_{p,\mathfrak{r}}(x,y)=f_{p,\mathfrak{r}^{\prime}}(x,y)$ and $s_{p,\mathfrak{r}}(x)=s_{p,\mathfrak{r}^{\prime}}(x)$ for all $p\in\mathfrak{P}_{k}$.
Let $c=0$ and $\theta$ be a complex number. Suppose $c_{F}^{\nu}$ is a complex number for all $\nu\in P^{+}$ and $F_{N}(x_{1},\ldots,x_{N})$ has $c_{F_{N}}^{\nu}=Nc_{F}^{\nu}$ for $\nu\in P$ with $1\leq\ell(\nu)\leq N$ and $c_{F_{N}}^{\nu}=0$ for $\nu\in P$ with $\ell(\nu)\geq N+1$. Then, from Lemma 3.11, Proposition 3.12, and Theorem 3.2 with $c=0$,
@@
where the $c_{F}^{\nu}$ and thus $p\in\mathfrak{P}_{k}$ are constants. Since $f_{p,\mathfrak{r}^{\prime}}(x,y)$ has degree at most $k-\ell(p)$ in $y$ from Proposition 3.10, $\lim_{N\rightarrow\infty}\frac{f_{p,\mathfrak{r}^{\prime}}(\theta,N)p}{N^{k-\ell(p)}}=s_{p,\mathfrak{r}^{\prime}}(\theta)p=s_{p,\mathfrak{r}}(\theta)p$. Then,
@@
Proof of Theorem 1.9.
First, we prove the analogue of Theorem 3.2 in this setting. To show that this implies the result, we can follow the argument in Section 4.2. Suppose $\lambda\in P^{+}$ and replace $\ln(G_{\theta_{N}}(x_{1},\ldots,x_{N};\mu_{N}))$ with $\tilde{F}_{N}(x_{1},\ldots,x_{N})$ as we do previously, where $\tilde{F}_{N}$ also depends on $|\lambda|$. Then, for $\nu\in P^{+}$ with $|\nu|\leq|\lambda|$,
@@
for any positive integers $i_{1},\ldots,i_{r}\leq N$ such that $\sigma((i_{1},\ldots,i_{r}))=\nu$, there exists $c_{\nu}\in\mathbb{R}_{\geq 0}$ such that $|c_{\nu}^{N}|\leq c_{\nu}+o_{N}(1)$ for all $\nu\in P^{+}$, and the goal is to show that
@@
Recall that the coefficient of $p$ in $[1]\mathcal{D}_{\mathfrak{r}}^{\theta_{N}}(\tilde{F}_{N}(x_{1},\ldots,x_{N}))$ is $f_{p,\mathfrak{r}}(\theta_{N},N)$. Consider the term $f_{p,\mathfrak{r}}(\theta_{N},N)p$. We have that $|p|=O(N^{(1-c)m})$. Let $g_{p,\mathfrak{r}}(\theta,N)$ denote the sum of the summands of $f_{p,\mathfrak{r}}(\theta,N)$ with degree $k-m$ in $\theta$; note that this is the maximal degree of $\theta$ because any sequence that contributes to the coefficient of $p$ must have exactly $m$ term multiplications. Because $|T_{\ell}|=O(N^{\ell})$ by the proof of Proposition 3.10,
@@
$|f_{p,\mathfrak{r}}-g_{p,\mathfrak{r}}|(\theta_{N},N)|p|=O(N^{(k-1)(1-c)}).$
@@
$\left|\sum_{p\in\mathfrak{P}_{k}^{N}}(f_{p,\mathfrak{r}}-g_{p,\mathfrak{r}})(\theta_{N},N)\right|=O(N^{(k-1)(1-c)}).$
@@
$[1]\mathcal{D}_{\mathfrak{r}}^{\theta_{N}}(\tilde{F}_{N}(x_{1},\ldots,x_{N}))=\sum_{p\in\mathfrak{P}_{k}^{N}}g_{p,\mathfrak{r}}(\theta_{N},N)p+O(N^{(k-1)(1-c)}).$
Let $h_{p,\mathfrak{r}}(\theta,N)$ be the term of $g_{p,\mathfrak{r}}(\theta,N)$ with degree $k-m$ in $N$; we have previously noted that $k-m$ is the maximal degree of $N$. Then, $h_{p,\mathfrak{r}}(\theta,N)$ is $(\theta N)^{k-m}$ multiplied by a rational number, which is possibly zero. Furthermore,
$[1]\mathcal{D}_{\mathfrak{r}}^{\theta_{N}}(\tilde{F}_{N}(x_{1},\ldots,x_{N}))=\sum_{p\in\mathfrak{P}_{k}^{N}}h_{p,\mathfrak{r}}(\theta_{N},N)p+O(N^{k(1-c)-1})+O(N^{(k-1)(1-c)}).$
From Lemma 6.3, $s_{p,\mathfrak{r}}(x)$ is a multiple of $x^{k-m}$. That is, the coefficient of $N^{k-m}$ in $f_{p,\mathfrak{r}}(\theta,N)$ is a multiple of $\theta^{k-m}$. This makes sense because the only sequences that can contribute to the coefficient $N^{k-m}$ must consist of $k-m$ switches, each of which contributes a factor of $\theta$. This implies that $h_{p,\mathfrak{r}}(\theta,N)=s_{p,\mathfrak{r}}(\theta)N^{k-m}$. Hence,
$[1]\mathcal{D}_{\mathfrak{r}}^{\theta_{N}}(\tilde{F}_{N}(x_{1},\ldots,x_{N}))=\sum_{p\in\mathfrak{P}_{k}^{N}}s_{p,\mathfrak{r}}(\theta)N^{k-m}p+O(N^{k(1-c)-1})+O(N^{(k-1)(1-c)}).$
@@
$=\sum_{\mathfrak{r}\in I_{N}(\lambda)}\sum_{p\in\mathfrak{P}_{k}^{N}}s_{p,\mathfrak{r}}(\theta_{N})N^{k-\ell(p)}p+O(N^{k(1-c)-1})+O(N^{(k-1)(1-c)})$
$=\sum_{p\in\mathfrak{P}_{k}^{N}}s_{\lambda,p}(\theta_{N})N^{k-\ell(p)+\ell(\lambda)}p+O(N^{\ell(\lambda)+k(1-c)-1})+O(N^{\ell(\lambda)+(k-1)(1-c)}).$
We explain how we arrive at this expression. Similarly to the proof of Proposition 3.12, for $\mathfrak{r}\in I_{N}(\lambda)$ such that $\sigma(\mathfrak{r})\not=\lambda$, for all $p\in\mathfrak{P}_{k}^{N}$ we have that
$s_{p,\mathfrak{r}}(\theta_{N})N^{k-\ell(p)}p=O(|\theta_{N}|^{k-\ell(p)}N^{k-\ell(p)})p=O(N^{k(1-c)}).$
Since the number of such $\mathfrak{r}$ is $O(N^{\ell(\lambda)-1})$, this contributes to the $O(N^{\ell(\lambda)+k(1-c)-1})$ remainder term.
Next, observe that $s_{\lambda,p}(\theta_{N})=(1+o_{N}(1))s_{\lambda,p}(\theta)N^{-c(k-\ell(p))}$. Therefore,
$\sum_{p\in\mathfrak{P}_{k}^{N}}s_{\lambda,p}(\theta_{N})N^{k-\ell(p)+\ell(\lambda)}$ $=(1+o_{N}(1))\sum_{p\in\mathfrak{P}_{k}^{N}}s_{\lambda,p}(\theta)N^{-c(k-\ell(p))}N^{k-\ell(p)+\ell(\lambda)}p$
$=(1+o_{N}(1))N^{\ell(\lambda)}\sum_{p\in\mathfrak{P}_{k}^{N}}s_{\lambda,p}(\theta)N^{(1-c)(k-\ell(p))}p$
$=(1+o_{N}(1))N^{\ell(\lambda)+k(1-c)}\sum_{p\in\mathfrak{P}_{k}^{N}}s_{\lambda,p}(\theta)\frac{p}{N^{\ell(p)(1-c)}}.$
@@
$=(1+o_{N}(1))N^{\ell(\lambda)+k(1-c)}\sum_{p\in\mathfrak{P}_{k}^{N}}s_{\lambda,p}(\theta)\frac{p}{N^{\ell(p)(1-c)}}+O(N^{\ell(\lambda)+k(1-c)-1})+O(N^{\ell(\lambda)+(k-1)(1-c)}).$
Afterwards, Lemma 6.3 with $\lambda$ as $\mathfrak{r}$ gives that
(24) $\begin{split}&\frac{1}{N^{\ell(\lambda)+k(1-c)}}\sum_{\mathfrak{r}\in I_{N}(\lambda)}[1]\mathcal{D}_{\mathfrak{r}}^{\theta_{N}}(\tilde{F}_{N}(x_{1},\ldots,x_{N}))\\ &=(1+o_{N}(1))\sum_{p\in\mathfrak{P}_{k}^{N}}s_{\lambda,p}(\theta)\frac{p}{N^{\ell(p)(1-c)}}+O(N^{-1})+O(N^{-(1-c)})\\ &=(1+o_{N}(1))\prod_{i=1}^{\ell(\lambda)}\left(\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}\theta^{|B|-1}\left(\sum_{\nu\in P,|\nu|=|B|}(-1)^{\ell(\nu)-1}\frac{|\nu|P(\nu)}{\ell(\nu)}c_{N}^{\nu}\right)\right)+o_{N}(1),\end{split}$
@@
Afterwards, we can show (3) to prove the forward direction using induction on $k$.
@@
$[1]\mathcal{D}_{\mathfrak{r}}^{\theta}(F(x_{1},\ldots,x_{N}))=\sum_{p\in\mathfrak{P}_{k}}(s_{p,\mathfrak{r}}(\theta)N^{k-\ell(p)}+R_{1,\mathfrak{r},p}(\theta,N)+R_{2,\mathfrak{r},p}(\theta,N))p,$
@@
  • The maximum degree of $\theta$ in the polynomial $R_{1,\mathfrak{r},p}(\theta,N)$ is at most $k-\ell(p)-1$ and the maximum degree of $N$ in the coefficient of $\theta^{\ell}$ is at most $\ell$ for $10\leq\ell\leq k-\ell(p)-$ 1.
@@
is the limit of the coefficient of $p$ in [matrix]*(5.2) as $N\rightarrow\infty$, where $\alpha_{i}$ is the coefficient of $\theta^{i}N^{i}$ in $R_{1,\mathfrak{r},p}(\theta,N)$ for $10\leq i\leq k-\ell(p)-$ 1.
Furthermore, we have the following corollary, which is analogous to Corollary 5.4.
@@
Assume that for all $\nu\in P^{+}$, there exists $c_{\nu}\in\mathbb{R}_{\geq 0}$ such that $|c_{\nu}^{N}|\leq c_{\nu}+o_{N}(1)$. Then, $\{\mu_{N}\}_{N\geq 1}$ satisfies a $(c+\Delta)$-LLN with free cumulants $\{c_{k}\}_{k\geq 1}$ if and only if
@@
We follow the method of the proof of Corollary 5.4. For $D=N^{\Delta}$, let $\mu_{N}^{D}$ be the pushforward of $\mu_{N}$ with respect to the function $f:A\rightarrow A,x\mapsto Dx$. By Lemma 5.2, for all $\nu\in P^{+}$,
@@
Since $\Delta|\nu|-\alpha_{\nu}\leq-c$, it is clear that $\left|\frac{N^{\Delta|\nu|-\alpha_{\nu}}}{N^{-c}}c_{\nu}^{N}\right|=O_{N,\nu}(1)$ for $\nu\in P^{+}$. Then, by Theorem 1.9, $\{\mu_{N}^{D}\}_{N\geq 1}$ satisfy a $c$-LLN with free cumulants $\{c_{k}\}_{k\geq 1}$ if and only if

### v3

@@ 6. Polynomial Coefficients
When $1]\mathcal{D}^{\theta}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))$ is expressed as a polynomial in the $c_{F}^{\nu}$ for $\nu\in P^{+}$, we expect the coefficients to be polynomials in $\theta$ and $N$. In this section, we show that this is the case and relate the leading terms of the polynomial coefficients with the main results. In fact, we use this characterization of the coefficients to prove Theorem˜1.8, which generalizes Theorem˜1.6.
6.1. Proof of Theorem˜1.8
First, we state the following result, which is a straightforward implication of Lemma˜4.8.
@@
    If $\sigma(\mathfrak{r}_{1})=\sigma(\mathfrak{r}_{2})$, then $1]\mathcal{D}_{\mathfrak{r}_{2}}^{\theta}(F(x_{1},\ldots,x_{N}))=[1]\mathcal{D}_{\mathfrak{r}_{1}}^{\theta}(F(x_{1},\ldots,x_{N}))$ for $N\geq\max(\mathfrak{r}_{1},\mathfrak{r}_{2})$.
@@
Suppose $\mathfrak{r}_{1}=\{i_{j}^{1}\}_{1\leq j\leq k}$ and $\mathfrak{r}_{2}=\{i_{j}^{2}\}_{1\leq j\leq k}$, where $k\geq 1$. Clearly, there is nothing to prove if $\mathfrak{r}_{1}$ and $\mathfrak{r}_{2}$ have different lengths. Statement (a) follows from Lemma˜4.8, because
@@
there exists a polynomial $f(x,y)$ with rational coefficients such that for all complex numbers $\theta$ and $N\geq\max(\mathfrak{r})$, the coefficient of $p$ in $1]\mathcal{D}^{\theta}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))$ is $f(\theta,N)$.
@@
[…]
If the sum of the $|\nu_{i}|$ is not $k$, from Corollary˜3.6 the coefficient of $p$ is 0. Suppose that the sum of the $|\nu_{i}|$ is $k$. Let $S$ be the set of elements in $\mathfrak{r}$. Also, for $s\in T_{N}$, let $X(s)$ be the set $x_{i}$ such that $i\notin\mathfrak{r}$ and $x_{i}$ appears in switches or term multiplications of $s$. Since each $x_{i}\in X(s)$ from term multiplications must be eliminated by at least one switch by ˜3.10.1, and each other $x_{i}\in X(s)$ appears in at least one switch, $|X(s)|\leq k$ because there are at most $k$ switches.
[…]
@@
[…]
@@
Using Lemma˜6.2, for indices $\mathfrak{r}$ of length $k$ and $N\geq\max(\mathfrak{r})$, let the coefficient of $p$ in $1]\mathcal{D}^{\theta}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))$ be $f_{p,\mathfrak{r}}(\theta,N)$ for $p$ in $\mathfrak{P}_{k}$. Then, for $N\geq\max(\mathfrak{r})$,
(23) $1]\mathcal{D}^{\theta}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))=\sum_{p\in\mathfrak{P}_{k}}f_{p,\mathfrak{r}}(\theta,N)p.$
Also, from Proposition˜3.10 with $c=0$ and $\theta_{N}=\theta$ for $N\geq 1$ for any $\theta\in\mathbb{C}$, $f_{p,\mathfrak{r}}(x,y)$ has degree at most $k-\ell(p)$ in $y$. Then, let the coefficient of $y^{k-\ell(p)}$ in $f_{p,\mathfrak{r}}(x,y)$ be $s_{p,\mathfrak{r}}(x)$, with $s_{p,\mathfrak{r}}(x)=0$ being possible.
@@
Proof of Lemma˜6.3.
Consider the indices $\mathfrak{r}^{\prime}$ of length $k$, where the first $\lambda_{1}$ indices are $1$, the next $\lambda_{2}$ are $2$, and so forth, until the last $\lambda_{m}$ indices are $m$. From Corollary˜6.1, $1]\mathcal{D}^{\theta}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))=[1]\mathcal{D}^{\theta}_{\mathfrak{r}^{\prime}}(F(x_{1},\ldots,x_{N}))$ for all $N\geq\max(\mathfrak{r})$. Therefore, $f_{p,\mathfrak{r}}(x,y)=f_{p,\mathfrak{r}^{\prime}}(x,y)$ and $s_{p,\mathfrak{r}}(x)=s_{p,\mathfrak{r}^{\prime}}(x)$ for all $p\in\mathfrak{P}_{k}$.
Let $c=0$ and $\theta$ be a complex number. Suppose $c_{F}^{\nu}$ is a complex number for all $\nu\in P^{+}$ and $F_{N}(x_{1},\ldots,x_{N})$ has $c_{F_{N}}^{\nu}=Nc_{F}^{\nu}$ for $\nu\in P$ with $1\leq\ell(\nu)\leq N$ and $c_{F_{N}}^{\nu}=0$ for $\nu\in P$ with $\ell(\nu)\geq N+1$. Then, from Lemma˜3.11, Proposition˜3.12, and Theorem˜3.2 with $c=0$,
@@
where the $c_{F}^{\nu}$ and thus $p\in\mathfrak{P}_{k}$ are constants. Since $f_{p,\mathfrak{r}^{\prime}}(x,y)$ has degree at most $k-\ell(p)$ in $y$ from Proposition˜3.10, $\lim_{N\rightarrow\infty}\frac{f_{p,\mathfrak{r}^{\prime}}(\theta,N)p}{N^{k-\ell(p)}}=s_{p,\mathfrak{r}^{\prime}}(\theta)p=s_{p,\mathfrak{r}}(\theta)p$. Then,
@@
Proof of Theorem˜1.8.
First, we prove the analogue of Theorem˜3.2 in this setting. To show that this implies the result, we can follow the argument in Section˜4.2. Suppose $\lambda\in P^{+}$ and replace $\ln(G_{\theta_{N}}(x_{1},\ldots,x_{N};\mu_{N}))$ with $\tilde{F}_{N}(x_{1},\ldots,x_{N})$ as we do previously, where $\tilde{F}_{N}$ also depends on $|\lambda|$. Then, for $\nu\in P^{+}$ with $|\nu|\leq|\lambda|$,
@@
for any positive integers $i_{1},\ldots,i_{r}\leq N$ such that $\sigma((i_{1},\ldots,i_{r}))=\nu$, $|c_{\nu}^{N}|=N^{o_{N}(1)}$ for all $\nu\in P^{+}$, and the goal is to show that
@@
Recall that the coefficient of $p$ in $1]\mathcal{D}_{\mathfrak{r}}^{\theta_{N}}(\tilde{F}_{N}(x_{1},\ldots,x_{N}))$ is $f_{p,\mathfrak{r}}(\theta_{N},N)$. Consider the term $f_{p,\mathfrak{r}}(\theta_{N},N)p$. We have that $|p|=O(N^{(1-c)m+o_{N}(1)})$ by the $|c_{\nu}^{N}|=N^{o_{N}(1)}$ condition. Let $g_{p,\mathfrak{r}}(\theta,N)$ denote the sum of the summands of $f_{p,\mathfrak{r}}(\theta,N)$ with degree kmk-m in $\theta$; note that this is the maximal degree of $\theta$ because any sequence that contributes to the coefficient of $p$ must have exactly $m$ term multiplications. The values $T_{\ell}$ for $0\leq\ell\leq k-m-1$ are defined in the proof of Proposition˜3.10; the proof shows that $|T_{\ell}|=O(N^{\ell})$, so
@@
$|f_{p,\mathfrak{r}}-g_{p,\mathfrak{r}}|(\theta_{N},N)|p|=O(N^{(k-1)(1-c)+o_{N}(1)}).$
@@
$\left|\sum_{p\in\mathfrak{P}_{k}^{N}}(f_{p,\mathfrak{r}}-g_{p,\mathfrak{r}})(\theta_{N},N)\right|=O(N^{(k-1)(1-c)+o_{N}(1)}).$
@@
$1]\mathcal{D}_{\mathfrak{r}}^{\theta_{N}}(\tilde{F}_{N}(x_{1},\ldots,x_{N}))=\sum_{p\in\mathfrak{P}_{k}^{N}}g_{p,\mathfrak{r}}(\theta_{N},N)p+O(N^{(k-1)(1-c)+o_{N}(1)}).$
Let $h_{p,\mathfrak{r}}(\theta,N)$ be the term of $g_{p,\mathfrak{r}}(\theta,N)$ with degree kmk-m in $N$; we have previously noted that kmk-m is the maximal degree of $N$. Then, $h_{p,\mathfrak{r}}(\theta,N)$ is $(\theta N)^{k-m}$ multiplied by a rational number, which is possibly zero. Furthermore,
$1]\mathcal{D}_{\mathfrak{r}}^{\theta_{N}}(\tilde{F}_{N}(x_{1},\ldots,x_{N}))=\sum_{p\in\mathfrak{P}_{k}^{N}}h_{p,\mathfrak{r}}(\theta_{N},N)p+O(N^{k(1-c)-1+o_{N}(1)})+O(N^{(k-1)(1-c)+o_{N}(1)}).$
From Lemma˜6.3, $s_{p,\mathfrak{r}}(x)$ is a multiple of $x^{k-m}$. That is, the coefficient of $N^{k-m}$ in $f_{p,\mathfrak{r}}(\theta,N)$ is a multiple of $\theta^{k-m}$. This makes sense because the only sequences that can contribute to the coefficient $N^{k-m}$ must consist of kmk-m switches, each of which contributes a factor of $\theta$. This implies that $h_{p,\mathfrak{r}}(\theta,N)=s_{p,\mathfrak{r}}(\theta)N^{k-m}$. Hence,
$1]\mathcal{D}_{\mathfrak{r}}^{\theta_{N}}(\tilde{F}_{N}(x_{1},\ldots,x_{N}))=\sum_{p\in\mathfrak{P}_{k}^{N}}s_{p,\mathfrak{r}}(\theta_{N})N^{k-m}p+O(N^{k(1-c)-1+o_{N}(1)})+O(N^{(k-1)(1-c)+o_{N}(1)}).$
@@
(24) $\begin{split}&\sum_{\mathfrak{r}\in I_{N}(\lambda)}[1]\mathcal{D}_{\mathfrak{r}}^{\theta_{N}}(\tilde{F}_{N}(x_{1},\ldots,x_{N}))\\ &=\sum_{\mathfrak{r}\in I_{N}(\lambda)}\sum_{p\in\mathfrak{P}_{k}^{N}}s_{p,\mathfrak{r}}(\theta_{N})N^{k-\ell(p)}p+O(N^{k(1-c)-1})+O(N^{(k-1)(1-c)})\\ &=\sum_{p\in\mathfrak{P}_{k}^{N}}s_{p,\lambda}(\theta_{N})N^{k-\ell(p)+\ell(\lambda)}p+O(N^{\ell(\lambda)+k(1-c)-1+o_{N}(1)})+O(N^{\ell(\lambda)+(k-1)(1-c)+o_{N}(1)}).\end{split}$
We explain how we arrive at this expression. Similarly to the proof of Proposition˜3.12, for $\mathfrak{r}\in I_{N}(\lambda)$ such that $\sigma(\mathfrak{r})\not=\lambda$, for all $p\in\mathfrak{P}_{k}^{N}$ we have that
$s_{p,\mathfrak{r}}(\theta_{N})N^{k-\ell(p)}p=O(|\theta_{N}|^{k-\ell(p)}N^{k-\ell(p)})p=O(N^{k(1-c)+o_{N}(1)}).$
Since the number of such $\mathfrak{r}$ is $O(N^{\ell(\lambda)-1})$, this contributes to the $O(N^{\ell(\lambda)+k(1-c)-1+o_{N}(1)})$ remainder term.
Next, observe that $s_{p,\lambda}(\theta_{N})=(1+o_{N}(1))s_{p,\lambda}(\theta)N^{-c(k-\ell(p))}$. Therefore,
$\sum_{p\in\mathfrak{P}_{k}^{N}}s_{p,\lambda}(\theta_{N})N^{k-\ell(p)+\ell(\lambda)}p$ $=(1+o_{N}(1))\sum_{p\in\mathfrak{P}_{k}^{N}}s_{p,\lambda}(\theta)N^{-c(k-\ell(p))}N^{k-\ell(p)+\ell(\lambda)}p$
$=(1+o_{N}(1))N^{\ell(\lambda)}\sum_{p\in\mathfrak{P}_{k}^{N}}s_{p,\lambda}(\theta)N^{(1-c)(k-\ell(p))}p$
$=(1+o_{N}(1))N^{\ell(\lambda)+k(1-c)}\sum_{p\in\mathfrak{P}_{k}^{N}}s_{p,\lambda}(\theta)\frac{p}{N^{\ell(p)(1-c)}}.$
@@
$\sum_{\mathfrak{r}\in I_{N}(\lambda)}[1]\mathcal{D}_{\mathfrak{r}}^{\theta_{N}}(\tilde{F}_{N}(x_{1},\ldots,x_{N}))=$
$(1+o_{N}(1))N^{\ell(\lambda)+k(1-c)}\sum_{p\in\mathfrak{P}_{k}^{N}}s_{p,\lambda}(\theta)\frac{p}{N^{\ell(p)(1-c)}}+O(N^{-1+o_{N}(1)})+O(N^{-(1-c)+o_{N}(1)}).$
Afterwards, Lemma˜6.3 with $\lambda$ as $\mathfrak{r}$ gives that
(25) $\begin{split}&\frac{1}{N^{\ell(\lambda)+k(1-c)}}\sum_{\mathfrak{r}\in I_{N}(\lambda)}[1]\mathcal{D}_{\mathfrak{r}}^{\theta_{N}}(\tilde{F}_{N}(x_{1},\ldots,x_{N}))\\ &=(1+o_{N}(1))\sum_{p\in\mathfrak{P}_{k}^{N}}s_{p,\lambda}(\theta)\frac{p}{N^{\ell(p)(1-c)}}+O(N^{-1+o_{N}(1)})+O(N^{-(1-c)+o_{N}(1)})\\ &=(1+o_{N}(1))\prod_{i=1}^{\ell(\lambda)}\left(\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}\theta^{|B|-1}\left(\sum_{\nu\in P,|\nu|=|B|}(-1)^{\ell(\nu)-1}\frac{|\nu|P(\nu)}{\ell(\nu)}c_{N}^{\nu}\right)\right)+o_{N}(1),\end{split}$
@@
Afterwards, we can show (2) using induction on $k$ to prove the forward direction.
@@
$1]\mathcal{D}_{\mathfrak{r}}^{\theta}(F(x_{1},\ldots,x_{N}))=\sum_{p\in\mathfrak{P}_{k}}(s_{p,\mathfrak{r}}(\theta)N^{k-\ell(p)}+R_{1,\mathfrak{r},p}(\theta,N)+R_{2,\mathfrak{r},p}(\theta,N))p,$
@@
  • The maximum degree of $\theta$ in the polynomial $R_{1,\mathfrak{r},p}(\theta,N)$ is at most $k-\ell(p)-1$ and the maximum degree of $N$ in the coefficient of $\theta^{\ell}$ is at most $\ell$ for $0\leq\ell\leq k-\ell(p)-1$.
@@
is the limit of the coefficient of $p$ in [matrix]*(5.2) as $N\rightarrow\infty$, where $\alpha_{i}$ is the coefficient of $\theta^{i}N^{i}$ in $R_{1,\mathfrak{r},p}(\theta,N)$ for $0\leq i\leq k-\ell(p)-1$.
Furthermore, we have the following corollary, which is analogous to Corollary˜5.4.
@@
Assume that for all $\nu\in P^{+}$, $|c_{\nu}^{N}|=N^{o_{N}(1)}$. Then, $\{\mu_{N}\}_{N\geq 1}$ satisfies a $(c+\Delta)$-LLN with free cumulants $\{c_{k}\}_{k\geq 1}$ if and only if
@@
We follow the method of the proof of Corollary˜5.4. For $D=N^{\Delta}$, let $\mu_{N}^{D}$ be the pushforward of $\mu_{N}$ with respect to the function $f:A\rightarrow A,x\mapsto Dx$. By Lemma˜5.2, for all $\nu\in P^{+}$,
@@
Since $\Delta|\nu|-\alpha_{\nu}\leq-c$, it is clear that $\left|\frac{N^{\Delta|\nu|-\alpha_{\nu}}}{N^{-c}}c_{\nu}^{N}\right|=N^{o_{N}(1)}$ for $\nu\in P^{+}$. Then, by Theorem˜1.8, $\{\mu_{N}^{D}\}_{N\geq 1}$ satisfy a $c$-LLN with free cumulants $\{c_{k}\}_{k\geq 1}$ if and only if